If the mean GPA among students is 3.25 with a standard deviation of 0.75, and we select a random sample of 300 people, at what value for the sample mean would be greater than exactly 95% of possible sample means?
Answer #1

For sampling distribution of mean, P(X < A) = P(Z < (A - mean)/standard error)

Sample size, n = 300

Mean = 3.25

Standard error = sigma / sqrt{n}

= 0.75/V300

= 0.0433

Let the sample mean that would be greater than exactly 95% of possible sample means be M

P(X < M) = 0.95

P(Z < (M - 3.25)/0.0433) = 0.95 (take the value of Z corresponding to 0.95 from standard normal distribution table)

(M - 3.25)/0.0433 = 1.645

M = 3.3212